Find the perimter and area of a 3 ft by 9 ft rectangle
Answer by ReadingBoosters(3246)

Perimeter is the distance around the rectangle: 3+9+3+9
Area is the amount within the rectangle:3*9
...
p = 2l+2w = 2(9) + 2(3) = 18 + 6 = 24 ft
...
a = l*w = 3*9 = 27 ft^2
..............................
Delighted to help; feel free to email directly at HomeworkHelpers@ReadingBoosters.com.
-Reading Boosters
Website: www.MyHomeworkAnswers.com